A round bottom glass flask is filled to the mark with 250cm3 of mercury at 20°C. if the flask and its contents are heated to 40°C, how much mercury will be above the mark?
αglass = 9 x 10-6 °C-1
βmercury = 182 x 10-6 °C-1
1
Expert's answer
2020-11-05T10:32:36-0500

Coefficient of volume expansion of mercury βmercury = 182 x 10-6 °C-1

∆T = 40 – 20 = 20 °C

V0 = 250 cm3

Increase in volume of mercury:

"∆V = V_0\times β \times ∆T"

"∆V = 250 \times 182 \times 10^{-6} \times 20 = 0.91 \;cm^3"

Answer: 0.91 cm3